# Weil Restriction and the Motivic Cycle Class Map

### v1

@@ Abstract.
The étale cycle class map links Chow groups with étale cohomology. In the framework of motivic cohomology, we have the motivic cycle class map, which appears as a comparison morphism relating motivic and Lichtenbaum cohomology. By an examination of the comparison morphism constructed by Geisser and Levine, we show that the motivic cycle class map is induced by the étale cycle class map. We then study the behavior of cycle class maps under Weil restriction. Extending Karpenko’s construction of the Weil restriction for algebraic cycles and Chow groups, we introduce a corresponding Weil restriction for $\ell$-adic cohomology and prove its compatibility with the motivic cycle class map. This yields a conceptual explanation of the descent of cycle classes under finite Galois extensions.
@@
Weil restriction; Motivic cohomology; Étale realization; Chow rings. Mathematical Subject Classification (2020): 14F42; 14C25
@@
This naturally leads to the question of how different cohomology theories are related. As a bridge between such theories, the étale cycle class map provides an initial link between Chow theory—a Borel–Moore oriented cohomology theory in the sense of Levine and Morel [9]—and étale cohomology.
More concretely, for a smooth variety $X$ of finite type over a field $k$, there is a ring homomorphism
$cl_{X}^{*}\colon\mathop{\mathrm{CH}}\nolimits^{*}(X)\longrightarrow H^{2*}\!\left(X_{\acute{e}t},\mathbb{Q}_{\ell}(*)\right),$
under which the intersection product on $\mathop{\mathrm{CH}}\nolimits^{*}(X)$ corresponds to the cup product in étale cohomology. Here $\mathbb{Q}_{\ell}(*)$ denotes the $\ell$-adic coefficient system equipped with the appropriate Tate twist.
Another important construction in algebraic geometry is the Weil restriction, a powerful tool for descent with numerous applications in arithmetic and geometry. The Weil restriction for algebraic cycles—and for Chow groups—was established by Karpenko [7]. In parallel with Karpenko’s work, we construct the Weil restriction map for $\ell$-adic cohomology.
The relationship between Chow groups and étale cohomology can be further clarified within the framework of motivic cohomology—conjectured by Beilinson and Lichtenbaum and later constructed by Voevodsky. In this setting, the Chow theory and étale cohomology are connected naturally as a comparison morphism arising from the change of topology from the Nisnevich site to the étale site, relating motivic cohomology to Lichtenbaum cohomology, which is known as the motivic cycle class map.
In this paper, we show that the the motivic cycle class map is indeed induced by the étale cycle class map. This is achieved by a careful examination of the construction of the comparison morphism due to Geisser and Levine [3] using Bloch’s cycle complex.
We then investigate the compatibility between Chow groups and étale cohomology under Weil restriction. In particular, we study this compatibility within the framework of motivic cohomology.
@@ 1.1. Weil restriction of schemes
We briefly review basics of Weil restriction of schemes. A detailed discussion can be found in [6] or [1].
@@ 1.2. Weil restriction of algebraic cycles
Pullbacks and pushforwards are functorial with respect to composition. Moreover, they respect the relation of rational equivalence; therefore, both operations descend to homomorphisms on Chow groups (see [2, Chapter I]).
Now we can start the construction of Weil restriction of algebraic cycles. Let $L/k$ be a finite Galois extension and let $X$ be a smooth quasi-projective scheme over $L$. The Weil restriction $\mathcal{R}(X)$ exists and is a smooth quasi-projective scheme over $k$ [7, Corollary 1.4]. Let $G=\mathop{\mathrm{Gal}}(L/k)$ be the Galois group.
@@
The key point is that the image of $p^{*}$ is precisely the subgroup of $G$-invariant cycles in $Z(\mathcal{R}(X)_{L})$ (see [7, Lemma 2.1]). In other words, there is an isomorphism
@@
which we call the Weil restriction of algebraic cycles. The Weil restriction map is compatible with flat pullbacks, proper pushforwards, and both interior and exterior products of cycles [7, Proposition 2.5].
Moreover, the Weil restriction of algebraic cycles is compatible with the relation of rational equivalence [7, Corollary 3.3]. Hence, the map descends to the level of Chow groups,
@@
It remains compatible with pushforwards, pullbacks, products of cycles, and Gysin pullbacks on Chow groups. Finally, it is compatible with the intersection product [7, Proposition 3.4]. For a modern proof of compatibility with the intersection product, we refer the reader to [8], where Fulton’s deformation to the normal cone is used.
@@
We fix a base field $k$. Throughout this section, a sheaf is an étale sheaf; any cohomology group will mean étale cohomology, defined as sheaf cohomology on the étale site unless otherwise specified. Recall that for an abelian group $A$, we denote by $\underline{A}$ the associated constant étale sheaf.
@@
Let $X$ be a smooth variety over $k$, and let $\Lambda:=\mathbb{Z}/\ell\mathbb{Z}$ with $\ell\neq\mathrm{char}(k)$. For a smooth closed subvariety $i\colon Z\hookrightarrow X$ of codimension $r$, we have the Gysin map (see [11, §5, Chapter VI])
@@
For a possibly singular closed subvariety $Z^{\prime}\subset X$, one can still define such a class using semi-purity [5, Proposition 2.2.6], which yields an identification
@@
Finally, by extending the construction linearly, we obtain the $r$th étale cycle class map
@@
By taking the direct sum over all codimensions $r\geq 0$, we obtain the total étale cycle class map
@@
The étale cycle class map enjoys good compatibility with operations on algebraic cycles. We record some of these properties in the following theorem.
Theorem 1.6 ([11, §9, Chapter VI]).
Let $X$ be a smooth variety over $k$. Under the above assumptions, the following hold:
@@
    For any finite étale morphism $f\colon Y\to X$, the étale cycle class map is compatible with the pullback of algebraic cycles along $f$.
@@
    The étale cycle class map is compatible with exterior products of algebraic cycles via the Künneth map in cohomology.
@@
    For two algebraic cycles that intersect transversely, the étale cycle class map is compatible with the intersection product via the cup product in cohomology.
@@ 2.1. Comparison map
Fix a field $k$. Let $\textbf{Sm}_{k}$ denote the category of smooth schemes over $k$. For each integer $q$, there is a complex (the Voevodsky–Suslin complex) of Nisnevich sheaves with transfers, denoted $\mathbb{Z}(q)$, on $\textbf{Sm}_{k}$ (equipped with the Nisnevich topology). The motivic cohomology of a smooth scheme $X$ is defined through the hypercohomology $H^{p,q}(X,\mathbb{Z}):=\mathbb{H}^{p}_{\mathrm{Nis}}\!\bigl(X,\mathbb{Z}(q)\bigr)$. (Indeed, $\mathbb{Z}(q)$ is also a complex of Zariski sheaves. For a detailed discussion, see [10, Part 3])
@@
where $\underline{\Lambda}(q)$ denotes the $q$th Tate twist for $\underline{\Lambda}:=\underline{\mathbb{Z}/l\mathbb{Z}}$ (See Definition 1.3)
@@
Let $X$ be a smooth variety over $k$. Is the motivic cycle class map $cl_{\mathrm{mot}}^{q}$ induced by the étale cycle class map considered in §1.3?
@@ 2.2. Naturality of the motivic cycle class map
Let $X$ be a smooth scheme over $k$, and let $\Lambda:=\mathbb{Z}/\ell\mathbb{Z}$ with $\ell\neq\mathrm{char}(k)$. For any integer $q\geq 0$, let $z^{q}(X,*)$ denote Bloch’s cycle complex (of Zariski sheaves) computing higher Chow groups [10, Part 5]. It is known that, after a suitable reindexing, the complex $z^{q}(X,*)$ is quasi-isomorphic to the Voevodsky–Suslin complex $\mathbb{Z}(q)$ (see [3, Section 2.5]). We denote the reindexed complex by $Z^{q}(X,*)$. In particular, its degree-$2q$ term satisfies $Z^{q}(X,2q)=Z^{q}(X)$ , the group of codimension-$q$ algebraic cycles on $X$.
Let $\pi:X_{\text{ét}}\to X_{Zar}$ be the canonical morphism from the étale site to the Zariski site. Then the construction of Geisser and Levine yields a natural map ([3, Equation 3.21])
@@
[3, cf. Theorem 1.1] Under the same assumptions as above, the map on hypercohomology induced by (2.4) coincides with the motivic cycle class map considered in §2.1 (i.e., (2.2)).
@@ 2.3. Motivic versus étale cycle class map
Let $X$ be a smooth variety over $k$, and let $\Lambda:=\mathbb{Z}/\ell\mathbb{Z}$ with $\ell\neq\mathrm{char}(k)$. Then the motivic cycle class map (2.2) is obtained from the étale cycle class map by tensoring with $\Lambda$ and passing to the quotient by rational equivalence relation.
@@
As a complex of étale sheaves, $\pi^{*}\!\bigl(\underline{\Lambda}(q)\bigr)$ is quasi-isomorphic to $G^{*}\!\bigl(X,\underline{\Lambda}(q)\bigr)$, the Godement resolution of $\underline{\Lambda}(q)$ ([3, Example 3.3]). Hence the map $\eta$ is reduced to
@@
(See [3, §3.7]) which yields a map $cl^{q}:Z^{q}(X)\otimes\Lambda\to H^{2q}(G^{*}(X,\underline{\Lambda}(q)))$.
@@
From [3, Lemma 3.8], we obtain the following commutative diagram:
@@
Viewing $W$ as a cycle in both $Z^{q}_{W}(X)$ and $Z^{q}(X)$, we conclude that
@@ 3.1. Weil restriction on $\ell$-adic cohomology
In this section, we construct the Weil restriction on $\ell$-adic cohomology. Unless otherwise specified, all cohomology groups will refer to étale cohomology.
@@
There are two reasons to consider $\ell$-adic cohomology. First, the étale cycle class maps (1.4) and (1.5) do not, in general, extend to Chow groups without additional hypotheses on the dimension of the ambient scheme. In contrast, for $\ell$-adic cohomology, such an extension is always possible [11, Remark 10.8, §10, Chapter VI]. Second, $\ell$-adic cohomology has good descent properties, which are described in the following lemma.
@@
Note that the projection morphism $X_{L}\to X$ is a finite Galois covering morphism (In the sense of [11, Remark 5.4, §5, Chapter I]). Hence we have the Hoschild-Serre spectral sequence
@@
([11, Theorem 2.20,§2, Chapter III]) Since $H^{q}(X_{L},\mathbb{Q}_{l}(r))$ is a vector space over $\mathbb{Q}_{l}$, which is uniquely $n$-divisible, we have that $E^{p,q}_{2}=0$ except for the term $p=0$ ([4, Corollary 3.3.9, Chapter III]). Hence we have
@@
We now turn to the construction of the Weil restriction in $\ell$-adic cohomology. Let $L/k$ be a finite Galois extension with Galois group $G$, and let $X$ be a smooth quasi-projective variety over $L$. We assume that $[L:k]=n$. Still, we denote by $\mathcal{R}(X)$ the Weil restriction of $X$ along $L/k$.
@@
Since $X$ is quasi-projective, one can describe $f^{\sigma}$ explicitly using Čech cohomology (see [11, Theorem 2.17, §2, Chapter III]). Let $U\to X$ be an étale covering morphism, and set
@@
Consider the following Künneth product homomorphism ([11, §8, Chapter VI]):

### v2

@@ Abstract.
We construct the Weil restriction map for $\ell$-adic cohomology and, more generally, for mixed Weil cohomology theories. We study its compatibility with the motivic cycle class map and show that these constructions admit a natural interpretation in the triangulated categories of motives. Using Grothendieck’s six-functor formalism, we prove that the Weil restriction map arises intrinsically from the functorial structures of these categories. This provides a conceptual framework for understanding the interaction between Weil restriction, motivic cohomology, and realization functors.
@@
Motives; Weil restriction; Motivic cohomology; Étale realization; Chow rings. Mathematical Subject Classification (2020): 14F42; 14C25
@@
This naturally leads to the question of how different cohomology theories are related. As a bridge between such theories, the étale cycle class map provides an initial link between the Chow theory—a Borel–Moore oriented cohomology theory in the sense of Levine and Morel [11]—and étale cohomology. Moreover, this relationship can be further clarified within the framework of motivic cohomology—conjectured by Beilinson and Lichtenbaum and later constructed by Voevodsky. In this setting, the Chow theory and the étale cohomology are connected naturally as a comparison map arising from the canonical change of site morphism, relating motivic cohomology to Lichtenbaum cohomology. This comparison map is known as the motivic cycle class map.
In this paper, we study the motivic cycle class map from a descent point of view. This is achieved by considering an important construction in algebraic geometry—the Weil restriction—which provides a powerful tool for descent and has numerous applications in arithmetic and geometry.
The Weil restriction for algebraic cycles, and for Chow groups, was established by Karpenko [9]. In parallel with Karpenko’s work, we construct the Weil restriction map for a large class of cohomology theories with good descent properties, including $\ell$-adic cohomology.
By working in the triangulated categories that give rise to these cohomology theories, we observe that our construction is intrinsic, since it relies solely on Grothendieck’s six-functor formalism in these categories. As a result, this paper provides an attempt to understand how concrete constructions and phenomena in algebraic geometry are governed by the six-functor formalism.
We organize the paper as follows. Section §1 is mainly preliminary, in which we recall the construction and basic properties of the Weil restriction of schemes and algebraic cycles. In Section §2, we show that the motivic cycle class map, viewed as a comparison map, is induced by the étale cycle map. In Section §3, we construct the Weil restriction map for $\ell$-adic cohomology and prove a compatibility theorem. Finally, in Section §4, we generalize our construction and compatibility theorem to mixed Weil cohomology theories.
Acknowledgments
The second author thanks Prof. Stefan Gille for introducing the problem of Weil restriction. The second author also thanks Prof. Nikita Karpenko for helpful discussions concerning his work on algebraic cycles.
@@ Acknowledgments
An algebra is a commutative algebra and a module means a module over a commutative ring.
@@ 1.1. Weil restriction of schemes
We briefly review the basics of Weil restriction of schemes. A detailed discussion can be found in [8] or [1].
@@
  • (4)
    If $X$ is a (quasi-)projective scheme over $L$, then $\mathcal{R}(X)$ is also (quasi-)projective over $k$.
@@ 1.2. Weil restriction of algebraic cycles
Pullbacks and pushforwards are functorial with respect to composition. Moreover, they respect the relation of rational equivalence; therefore, both operations descend to homomorphisms on Chow groups (see [4, Chapter I]).
Now we can start the construction of Weil restriction of algebraic cycles. Let $L/k$ be a finite Galois extension and let $X$ be a smooth quasi-projective scheme over $L$. The Weil restriction $\mathcal{R}(X)$ exists and is a smooth quasi-projective scheme over $k$ [9, Corollary 1.4]. Let $G=\mathop{\mathrm{Gal}}(L/k)$ be the Galois group.
@@
The key point is that the image of $p^{*}$ is precisely the subgroup of $G$-invariant cycles in $Z(\mathcal{R}(X)_{L})$ (see [9, Lemma 2.1]). In other words, there is an isomorphism
@@
which we call the Weil restriction of algebraic cycles. The Weil restriction map is compatible with flat pullbacks, proper pushforwards, and both interior and exterior products of cycles [9, Proposition 2.5].
Moreover, the Weil restriction of algebraic cycles is compatible with the relation of rational equivalence [9, Corollary 3.3]. Hence, the map descends to the level of Chow groups,
@@
It remains compatible with pushforwards, pullbacks, products of cycles, and Gysin pullbacks on Chow groups. Finally, it is compatible with the intersection product [9, Proposition 3.4]. For a modern proof of compatibility of Weil restriction map with the intersection product, we refer the reader to [10], where Fulton’s deformation to the normal cone is used.
@@
We fix a base field $k$. Throughout this section, a sheaf is an étale sheaf; any cohomology group will mean étale cohomology group, defined as sheaf cohomology on the étale site unless otherwise specified. Recall that for an abelian group $A$, we denote by $\underline{A}$ the associated constant étale sheaf.
@@
Let $X$ be a smooth scheme over $k$, and let $\Lambda:=\mathbb{Z}/\ell\mathbb{Z}$ with $\ell\neq\mathrm{char}(k)$. For a smooth closed subvariety $i\colon Z\hookrightarrow X$ of codimension $r$, we have the Gysin map (see [13, §5, Chapter VI])
@@
For a possibly singular closed subvariety $Z^{\prime}\subset X$, one can still define such a class using semi-purity [7, Proposition 2.2.6], which yields an identification
@@
Finally, by extending the construction linearly, we obtain the $r$th étale cycle map
@@
By taking the direct sum over all codimensions $r\geq 0$, we obtain the total étale cycle map
@@
The étale cycle map enjoys good compatibility with operations on algebraic cycles. We record some of these properties in the following theorem.
Theorem 1.6 ([13, §9, Chapter VI]).
Let $X$ be a smooth scheme over $k$. Under the above assumptions, the following hold:
@@
    For any finite étale morphism $f\colon Y\to X$, the étale cycle map is compatible with the pullback of algebraic cycles along $f$.
@@
    The étale cycle map is compatible with exterior products of algebraic cycles via the Künneth map in cohomology.
@@
    For two algebraic cycles that intersect transversely, the étale cycle map is compatible with the intersection product via the cup product in cohomology.
@@
In certain situations, for example when the étale cohomology groups have coefficients in a field of characteristic zero, the étale cycle map extends (with the same image) to a group homomorphism from the Chow group to the étale cohomology group with good functorial properties.
In this case, we refer to such a map as the étale cycle class map.
@@ 2.1. Comparison map
Fix a field $k$. Let $\textbf{Sm}_{k}$ denote the category of smooth schemes over $k$. For each integer $q$, there is a complex (the Voevodsky–Suslin complex) of Nisnevich sheaves with transfers, denoted $\mathbb{Z}(q)$, on $\textbf{Sm}_{k}$ (equipped with the Nisnevich topology). The motivic cohomology of a smooth scheme $X$ is defined through the hypercohomology $H^{p,q}(X,\mathbb{Z}):=\mathbb{H}^{p}_{\mathrm{Nis}}\!\bigl(X,\mathbb{Z}(q)\bigr)$. (Indeed, $\mathbb{Z}(q)$ is also a complex of Zariski sheaves. For a detailed discussion, see [12, Part 3])
@@
where $\underline{\Lambda}(q)$ denotes the $q$th Tate twist for $\underline{\Lambda}:=\underline{\mathbb{Z}/l\mathbb{Z}}$ (See Definition 1.3).
@@
Let $X$ be a smooth scheme over $k$. Is the motivic cycle class map $cl_{\mathrm{mot}}^{q}$ induced by the étale cycle map considered in §1.3?
@@ 2.2. Naturality of the motivic cycle class map
Let $X$ be a smooth scheme over $k$, and let $\Lambda:=\mathbb{Z}/\ell\mathbb{Z}$ with $\ell\neq\mathrm{char}(k)$. For any integer $q\geq 0$, let $z^{q}(X,*)$ denote Bloch’s cycle complex (of Zariski sheaves) computing higher Chow groups [12, Part 5]. It is known that, after a suitable reindexing, the complex $z^{q}(X,*)$ is quasi-isomorphic to the Voevodsky–Suslin complex $\mathbb{Z}(q)$ (see [5, Section 2.5]). We denote the reindexed complex by $Z^{q}(X,*)$. In particular, its degree-$2q$ term satisfies $Z^{q}(X,2q)=Z^{q}(X)$ , the group of codimension-$q$ algebraic cycles on $X$.
Let $\pi:X_{\text{ét}}\to X_{Zar}$ be the canonical morphism from the étale site to the Zariski site. Then the construction of Geisser and Levine yields a natural map ([5, Equation 3.21])
@@
[5, cf. Theorem 1.1] Under the same assumptions as above, the map on hypercohomology induced by (2.4) coincides with the motivic cycle class map considered in §2.1 (i.e., (2.2)).
@@ 2.3. Motivic versus étale cycle class map
Let $X$ be a smooth scheme over $k$, and let $\Lambda:=\mathbb{Z}/\ell\mathbb{Z}$ with $\ell\neq\mathrm{char}(k)$. Then the motivic cycle class map (2.2) is obtained from the étale cycle map by tensoring with $\Lambda$ and passing to the quotient by rational equivalence relation.
@@
As a complex of étale sheaves, $\pi^{*}\!\bigl(\underline{\Lambda}(q)\bigr)$ is quasi-isomorphic to $G^{*}\!\bigl(X,\underline{\Lambda}(q)\bigr)$, the Godement resolution of $\underline{\Lambda}(q)$ ([5, Example 3.3]). Hence the map $\eta$ is reduced to
@@
(See [5, §3.7]) which yields a map $cl^{q}:Z^{q}(X)\otimes\Lambda\to H^{2q}(G^{*}(X,\underline{\Lambda}(q)))$.
@@
From [5, Lemma 3.8], we obtain the following commutative diagram:
@@
Viewing $W$ as a cycle in both $Z^{q}_{W}(X)$ and $Z^{q}(X)$, note that the upper part of the digram is nothing more than the étale cycle map modulo $\Lambda$, we conclude that
@@ 3.1. Weil restriction on $\ell$-adic cohomology
In this section, we construct the Weil restriction on $\ell$-adic cohomology. Unless otherwise specified, all cohomology groups will refer to étale cohomology groups.
@@
There are two reasons to consider $\ell$-adic cohomology. First, the étale cycle class maps (1.4) and (1.5) do not, in general, extend to Chow groups without additional hypotheses on the dimension of the ambient scheme. In contrast, for $\ell$-adic cohomology, such an extension is always possible [13, Remark 10.8, §10, Chapter VI]. Second, $\ell$-adic cohomology has good descent properties, which are described in the following lemma.
@@
Note that the projection morphism $X_{L}\to X$ is a finite Galois covering morphism (In the sense of [13, Remark 5.4, §5, Chapter I]). Hence we have the Hoschild-Serre spectral sequence
@@
([13, Theorem 2.20,§2, Chapter III]).
Since $H^{q}(X_{L},\mathbb{Q}_{l}(r))$ is a vector space over $\mathbb{Q}_{l}$, which is uniquely $n$-divisible, we have that $E^{p,q}_{2}=0$ except for the term $p=0$ ([6, Corollary 3.3.9, Chapter III]). Hence we have
@@
We now turn to the construction of the Weil restriction on $\ell$-adic cohomology. Let $L/k$ be a finite Galois extension with Galois group $G$, and let $X$ be a smooth quasi-projective scheme over $L$. We assume that $[L:k]=n$. Still, we denote by $\mathcal{R}(X)$ the Weil restriction of $X$ along $L/k$.
@@
Since $X$ is quasi-projective, one can describe $f^{\sigma}$ explicitly using Čech cohomology (see [13, Theorem 2.17, §2, Chapter III]). Let $U\to X$ be an étale covering morphism, and set
@@
Consider the following Künneth product homomorphism ([13, §8, Chapter VI]):